OpenStudy (anonymous):

find the slope intercept form of an equation of a line passing thru (2,3) and a) parallel to y= -3x+2 b) perpendicular to y=-3x+2 plz help,... im effin stuck someone smart

OpenStudy (anonymous):

parallel means slope is the same, so use \[y-y_1=m(x-x_1)\] with \[x_1=2, y_1=3, m =-3\]

OpenStudy (anonymous):

get \[y-3=-3(x-2)\] \[y-3=-3x+6\] \[y=-3x+9\] for the first one. second one is similar, but this time slope is \[\frac{1}{3}\]